OpenStudy (pphalke):

Malcolm paid an average of $31 per tie for a recent purchase of three ties. The price of one tie was twice as much as another, and the remaining tie cost $24, what were the prices of the other two ties? what was the price of the less expensive of the 2 remaining ties?

OpenStudy (mathstudent55):

Again?

OpenStudy (pphalke):

I didn't add that question where it said what was the price of the less expensive of the 2 remaining ties

OpenStudy (faiqraees):

First tie = 2 times the second tie Third tie = 24 (First tie+Second tie +third tie)/3 = 31 24+x+2x=31*3

OpenStudy (faiqraees):

The value of x is the inexpensive one

OpenStudy (mathstudent55):

You already know the given price is $24. You calculated one other price to be $23. The third tie is twice $23. What is twice $23?

OpenStudy (pphalke):

that's 23 times 2 its 46

OpenStudy (mathstudent55):

Good. Given price: $24 Two remaining ties: More expensive one: $46 Less expensive one: $23